# RMDM: Radio Map Diffusion Model with Physics Informed

### v2

@@
Abstract
With the rapid development of wireless communication technology, the efficient utilization of spectrum resources, optimization of communication quality, and intelligent communication have become critical. Radio map reconstruction is essential for enabling advanced applications, yet challenges such as complex signal propagation and sparse data hinder accurate reconstruction. To address these issues, we propose the Radio Map Diffusion Model (RMDM), a physics-informed framework that integrates Physics-Informed Neural Networks (PINNs) to incorporate constraints like the Helmholtz equation. RMDM employs a dual U-Net architecture: the first ensures physical consistency by minimizing PDE residuals, boundary conditions, and source constraints, while the second refines predictions via diffusion-based denoising. By leveraging physical laws, RMDM significantly enhances accuracy, robustness, and generalization. Experiments demonstrate that RMDM outperforms state-of-the-art methods, achieving NMSE of 0.0031 and RMSE of 0.0125 under the Static RM (SRM) setting, and NMSE of 0.0047 and RMSE of 0.0146 under the Dynamic RM (DRM) setting. These results establish a novel paradigm for integrating physics-informed and data-driven approaches in radio map reconstruction, particularly under sparse data conditions. The code can be found on the website: github.com/Hxxxz0/RMDM
1 Introduction
@@
With the continuous breakthroughs made by humans in information technology, some new concepts and technologies have begun to emerge, such as 5G, 6G communications, terahertz communications, and autonomous driving technology Jornet et al. (2024); Kadir et al. (2021); Yurtsever et al. (2020). In order to fully utilize the advantages of various technologies, It is particularly important to plan the base station layout and allocate frequency resources precisely and efficiently Webbink (1977).
@@
Figure 1: Schematic representation of the RMDM architecture integrating the Helmholtz equation for path loss computation. The dual UNet framework consists of UNet 1, embedding physical constraints via PINNs, and UNet 2, refining outputs through diffusion-based denoising. Guided by the Helmholtz equation and loss functions ($L_{\text{PDE}}$, $L_{\text{BC}}$, $L_{\text{Source}}$), the model ensures accurate and robust radio map predictions.
@@
Planning and allocation are closely linked to radio frequency (RF) radio maps. These maps are like a spatial snapshot of RF signals. They are really important because they show how signal power is spread out in different areas. This distribution, referred to as the geographical signal power spectral density (PSD), quantifies the accumulation of RF signal power as a function of spatial location and frequency. It also takes into account how these factors are related and how they might change over time. Geographical characteristics influence signal paths Di Matteo et al. (2021) based on RF propagation properties, with signal strength quantified in dBm and associated with power attenuation. Two principal mapping approaches Mantiply et al. (1997a) exist: field measurement, which uses professional RF equipment for precise power capture and attenuation assessment but yields localized, discontinuous data Mantiply et al. (1997b) and model prediction, such as ray-tracing methods, which employ mathematical and physical principles to analyze propagation environments comprehensively, enabling large-scale data generation without extensive on-site measurements Frei et al. (2009). However, manual collection methods have disadvantages such as high cost and low efficiency.
@@
Most current radiomap estimation approaches fall into two main categories: model-based and model-free methods, such as REM-GAN Zhang et al. (2023) and RadioGAT Li et al. (2024). Model-based methods rely on predefined signal propagation models, like the log-distance path loss model for Wi-Fi radio map reconstruction or interpolation based on thin-plate spline kernels Alonazi et al. (2017), but they struggle to capture complex environmental effects such as shadowing and obstacles. Model-free methods, on the other hand, explore neighborhood information without assuming specific models, employing techniques such as Radial Basis Function interpolation and inverse distance weighting. While more flexible, they depend heavily on the quality of observed samples and often assume uniform data distributions—an assumption that fails in practical scenarios like MDT, where user measurements are unevenly distributed. Additionally, variations in propagation models and parameters across training datasets further complicate the reconstruction process. As a result, integrating model-based and model-free methods offers a promising avenue for overcoming these challenges effectively Audiffren and Bresciani (2022).
@@
Diffusion models Nichol and Dhariwal (2021); Song et al. (2020) have emerged as a robust framework in generative modeling chen et al. (2024); Lai et al. (2024), achieving state-of-the-art results in image synthesis, super-resolution, and inpainting Yang et al. (2023). These models leverage iterative denoising to effectively model complex data distributions, enabling the generation of high-quality , photorealistic images. Their flexibility and mathematical rigor have facilitated advancements in diverse applications, including creative content generation and medical imaging Croitoru et al. (2023). Recently, diffusion models have been applied to radio map reconstruction, where capturing the spatial distribution of signal power is critical. For instance, RadioDiff Wang et al. (2024) demonstrates how these models address the challenges posed by irregular and sparse signal environments by simulating the propagation of radio waves Sizun (2005). However, the limitation of data volume is a major challenge faced by the entire model field L’heureux et al. (2017). It severely restricts the model’s ability to learn the laws of the complex physical world, and the diffusion model is no exception. After in-depth analysis, we decided to solve this problem by integrating physical laws.
@@
Just as humans use physical laws to better predict natural phenomena, do models benefit in the same way? Even physical phenomena that have not occurred or been observed can be predicted. To leverage this capability, we combined our model with the Physics-Informed Neural Network (PINN) Lawal et al. (2022), integrating physical constraints to enhance performance and improve generalization ability, surpassing traditional data-driven approaches. Unlike conventional neural networks that rely heavily on extensive datasets to identify patterns, PINNs incorporate physical principles, enabling the model to utilize prior knowledge and physical constraints. This hybrid approach effectively addresses challenges such as limited data availability and complex propagation environments. During PINN training Monaco and Apiletti (2023); Nabian et al. (2021), the optimization process minimizes not only the traditional data-fitting loss—representing the deviation between predicted and observed values but also the residuals of the embedded physical equations, ensuring the model adheres to underlying physical laws while improving predictive accuracy and robustness.
To effectively integrate these principles, we designed a dual U-Net architecture as part of our Radio Map Diffusion Model (RMDM). The first stage (U-Net 1) focuses on feature extraction and embedding physical information via PINNs, while the second stage (U-Net 2) Ronneberger et al. (2015) refines these features through a diffusion process to produce high-resolution radio map estimates. This two-stage framework incorporates mechanisms like physics-informed loss functions and boundary constraints to ensure consistency with the underlying physical laws. By leveraging the Helmholtz equation in the loss function design, we not only enhanced the model’s ability to generate accurate predictions but also improved its generalization capability in scenarios with sparse or unevenly distributed data.
@@
This integration of physical laws and machine learning principles enables RMDM to accurately capture complex spatial signal characteristics Raissi et al. (2019). In experiments conducted under the Static RM (SRM) setting, our model achieved an NMSE of 0.0031 and an RMSE of 0.0125, outperforming state-of-the-art (SOTA) approaches. Extensive experimental validation further highlights the model’s superior performance in reconstructing radio maps across various challenging scenarios, establishing a novel paradigm that combines physics-driven methodologies with data-driven techniques to address real-world signal propagation challenges Li et al. (2020).
@@
  • We propose the RMDM (Radio Map Diffusion Model), a physics-informed framework that integrates the Helmholtz equation and diffusion processes within a dual U-Net architecture. This design ensures physical consistency and significantly enhances the accuracy of radio map reconstruction in sparse and complex propagation scenarios.
@@
  • By incorporating Physics-Informed Neural Networks (PINNs), RMDM embeds physical laws into the learning process, enabling precise modeling of spatial signal characteristics, such as signal strength and path loss, while improving generalization across diverse environments.
@@
  • Extensive experimental validation demonstrates that RMDM achieves state-of-the-art performance, including an NMSE of 0.0031 and RMSE of 0.0125 under the Static RM (SRM) setting, highlighting its robustness and scalability for real-world signal propagation challenges.
2 Related Work
2.1 Radio Map Estimation
@@
Radio map estimation is essential for applications such as network planning and spectrum management. Traditional methods include model-based approaches, which utilize established signal propagation models like the log-distance path loss (LDPL) model for Wi-Fi radio map reconstruction Jung et al. (2011). Interpolation techniques, such as thin-plate splines Keller and Borkowski (2019), have also been employed to achieve spatially continuous estimations. However, these methods often assume ideal conditions and may not accurately capture complex environmental factors like shadowing and multipath effects, leading to limitations in heterogeneous settings.
2.2 Diffusion Models
Diffusion models have emerged as powerful generative frameworks capable of modeling complex data distributions. In the context of radio map estimation, models such as RadioDiff Nichol and Dhariwal (2021) and RM-Gen Luo et al. (2024) have been proposed. RadioDiff is an effective generative diffusion model for sampling-free dynamic radio map construction, while RM-Gen leverages conditional denoising diffusion probabilistic models to synthesize radio maps using minimal and readily collected data. These approaches address challenges associated with extensive data collection and computationally intensive simulations, offering cost-effective solutions for network optimization tasks.
2.3 Physics-Informed Neural Networks (PINNs)
Physics-Informed Neural Networks (PINNs) integrate physical laws, typically represented by partial differential equations (PDEs) Sloan et al. (2012), into the training process of neural networks. This integration ensures that the model’s predictions adhere to known physical principles, enhancing interpretability and generalization, especially in data-scarce scenarios. In the realm of radio map estimation, physics-inspired machine learning approaches have been introduced to reconstruct radio maps efficiently from sparse samples. By embedding the Helmholtz equation, which describes wave propagation phenomena, into the neural network’s loss function, the model can learn to predict radio signal behavior more accurately. This physics-informed approach mitigates the limitations of purely data-driven models, particularly in environments where data collection is challenging.
3 Method
@@
Figure 2: An illustration of the RMDM framework. (a) shows the pipeline overview, combining a condition model and a diffusion model for radio map reconstruction. Key components include (b) Physics-Informed Neural Networks (PINNs), which enforce physical constraints through PDEs, and (c) Radio Frequency Spatial Attention (RF-SA), which aligns spatial and frequency features to enhance accuracy. The framework leverages anchor conditions and matrix products to iteratively refine features, ensuring accurate and robust predictions in complex propagation scenarios.
@@
The radio map estimation task aims to predict the spatial distribution of wireless signal strength within a given area. Wireless signal propagation follows physical laws, such as electromagnetic wave equations and path loss models Jiang et al. (2024). Incorporating these physical constraints into learning-based models can significantly improve the accuracy and reliability of predictions.
@@
As illustrated in Figure 2, we propose a diffusion-based probabilistic framework that integrates a dual UNet architecture with a physics-informed neural network (PINN). The proposed framework is structured in two stages:
@@
  • Conditional Generation Model (UNet 1): This stage extracts input features and employs a PINN module to enforce adherence to physical constraints.
@@
3.1 Problem Formulation
Reconstructing the radio map involves modeling the spatial distribution of wireless signals, which can be formulated using the Helmholtz equation. This equation provides a physical foundation for capturing signal propagation dynamics:
$\nabla^{2}u(x,y)+k^{2}(x,y)u(x,y)=-f(x,y),\quad(x,y)\in\Omega.$ (1)
@@ 3.1 Problem Formulation
$\Delta_{h}u(i,j)$ $=\frac{u(i+1,j)+u(i-1,j)+u(i,j+1)+u(i,j-1)}{h^{2}}$ (2)
$\quad-\frac{4u(i,j)}{h^{2}},$
@@
$\mathcal{L}_{\text{PDE}}=\frac{1}{N_{\text{int}}}\sum_{(i,j)\in\Omega_{\text{int}}}[r(i,j)]^{2},$ (4)
@@
$\mathcal{L}_{\text{BC}}=\frac{1}{N_{\text{bc}}}\sum_{(i,j)\in\partial\Omega}[u(i,j)-u_{\text{BC}}(i,j)]^{2},$ (5)
@@
$\mathcal{L}_{\text{Source}}=\frac{1}{N_{\text{src}}}\sum_{(i,j)\in\Omega_{\text{src}}}[u(i,j)-u_{\text{src}}(i,j)]^{2},$ (6)
@@
$\mathcal{L}_{\text{PINN}}=\lambda_{\text{PDE}}\mathcal{L}_{\text{PDE}}+\lambda_{\text{BC}}\mathcal{L}_{\text{BC}}+\lambda_{\text{Source}}\mathcal{L}_{\text{Source}},$ (7)
@@
3.2 Model Architecture
Our model employs a dual U-Net architecture, where U-Net 1 is dedicated to feature extraction and the embedding of physical information, while U-Net 2 focuses on the denoising process during diffusion Kawar et al. (2022). To enhance feature transfer between the two U-Nets, we propose two innovative mechanisms: the Anchor Condition and RF-SA.
3.2.1 UNet 1: Conditional Generation Model with PINN Alignment

### v3

# Physics-Informed Representation Alignment for Sparse Radio-Map Reconstruction

@@
Abstract.
Radio map reconstruction is essential for enabling advanced applications, yet challenges such as complex signal propagation and sparse observational data hinder accurate reconstruction in practical scenarios. Existing methods often fail to align physical constraints with data-driven features, particularly under sparse measurement conditions. To address these issues, we propose Physics-Aligned Radio Map Diffusion Model (PhyRMDM), a novel framework that establishes cross-domain representation alignment between physical principles and neural network features through dual learning pathways. The proposed model integrates Physics-Informed Neural Networks (PINNs) with a representation alignment mechanism that explicitly enforces consistency between Helmholtz equation constraints and environmental propagation patterns. Experimental results demonstrate significant improvements over state-of-the-art methods, achieving NMSE of 0.0031 under Static Radio Map (SRM) conditions, and NMSE of 0.0047 with in Dynamic Radio Map (DRM) scenarios. The proposed representation alignment paradigm provides 37.2% accuracy enhancement in ultra-sparse cases (1% sampling rate), confirming its effectiveness in bridging physics-based modeling and deep learning for radio map reconstruction. The code can be found on the website: Code
Radio Map Generation, Generative Modeling, 6G Networks, Spatial Attention, Wireless Communications
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758161isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Machine learningccs: Computing methodologies Neural networksccs: Computing methodologies Probabilistic modelsccs: Information systems Wireless networks
1. Introduction
@@
With the continuous breakthroughs made by humans in information technology, some new concepts and technologies have begun to emerge, such as 5G, 6G communications, terahertz communications, and autonomous driving technology (Jornet et al., 2024; Kadir et al., 2021; Yurtsever et al., 2020). In order to fully utilize the advantages of various technologies, It is particularly important to plan the base station layout and allocate frequency resources precisely and efficiently (Webbink, 1977).
@@
Figure 1. Schematic of the PhyRMDM architecture integrating the Helmholtz equation for path loss. The dual UNet design uses UNet 1 with PINN-based physical constraints and UNet 2 for diffusion-based refinement. Guided by losses ($L_{\text{PDE}}$, $L_{\text{BC}}$, $L_{\text{Source}}$), it ensures accurate and robust radio map predictions.
@@
Planning and allocation are closely linked to radio frequency (RF) radio maps. These maps are like a spatial snapshot of RF signals. They are really important because they show how signal power is spread out in different areas. This distribution, referred to as the geographical signal power spectral density (PSD), quantifies the accumulation of RF signal power as a function of spatial location and frequency. It also takes into account how these factors are related and how they might change over time. Geographical characteristics influence signal paths (Di Matteo et al., 2021) based on RF propagation properties, with signal strength quantified in dBm and associated with power attenuation. Two principal mapping approaches (Mantiply et al., 1997) exist: field measurement, which uses professional RF equipment for precise power capture and attenuation assessment but yields localized, discontinuous data (Mantiply et al., 1997) and model prediction, such as ray-tracing methods, which employ mathematical and physical principles to analyze propagation environments comprehensively, enabling large-scale data generation without extensive on-site measurements (Frei et al., 2009). However, manual collection methods have disadvantages such as high cost and low efficiency.
@@
Most current radiomap estimation approaches fall into two main categories: model-based and model-free methods, such as REM-GAN (Zhang et al., 2023) and RadioGAT (Li et al., 2024). Model-based methods rely on predefined signal propagation models, like the log-distance path loss model for Wi-Fi radio map reconstruction or interpolation based on thin-plate spline kernels (Alonazi et al., 2017), but they struggle to capture complex environmental effects such as shadowing and obstacles. Model-free methods, on the other hand, explore neighborhood information without assuming specific models, employing techniques such as Radial Basis Function interpolation and inverse distance weighting. While more flexible, they depend heavily on the quality of observed samples and often assume uniform data distributions—an assumption that fails in practical scenarios like MDT, where user measurements are unevenly distributed. Additionally, variations in propagation models and parameters across training datasets further complicate the reconstruction process. As a result, integrating model-based and model-free methods offers a promising avenue for overcoming these challenges effectively (Audiffren and Bresciani, 2022).
@@
Diffusion models (Nichol and Dhariwal, 2021; Song et al., 2020) have emerged as a robust framework in generative modeling (chen et al., 2024; Lai et al., 2024), achieving state-of-the-art results in image synthesis (Rombach et al., 2022; Ramesh et al., 2021; Ning et al., 2024), inpainting (Yang et al., 2023) and other tasks (Hu et al., 2024; Song et al., 2022). These models leverage iterative denoising to effectively model complex data distributions, enabling the generation of high-quality (Ho et al., 2020; Ning et al., 2024), photorealistic images. Their flexibility and mathematical rigor have facilitated advancements in diverse applications, including creative content generation and medical imaging (Croitoru et al., 2023). Recently, diffusion models have been applied to radio map reconstruction, where capturing the spatial distribution of signal power is critical. For instance, RadioDiff (Wang et al., 2024) demonstrates how these models address the challenges posed by irregular and sparse signal environments by simulating the propagation of radio waves (Sizun, 2005). However, the limitation of data volume is a major challenge faced by the entire model field (L’heureux et al., 2017). It severely restricts the model’s ability to learn the laws of the complex physical world, and the diffusion model is no exception. After in-depth analysis, We decided to solve this problem by embedding physical laws to achieve the representational alignment of physical information.
@@
Just as humans use physical laws to better predict natural phenomena, do models benefit in the same way? Even physical phenomena that have not occurred or been observed can be predicted. To leverage this capability, we combined our model with the Physics-Informed Neural Network (PINN) (Lawal et al., 2022), integrating physical constraints to enhance performance and improve generalization ability, surpassing traditional data-driven approaches. Unlike conventional neural networks that rely heavily on extensive datasets to identify patterns, this way incorporate physical principles, enabling the model to utilize prior knowledge and physical constraints. This hybrid approach effectively addresses challenges such as limited data availability and complex propagation environments. During PINN training (Monaco and Apiletti, 2023; Nabian et al., 2021), the optimization process minimizes not only the traditional loss of data fit-representing not only the deviation between the predicted and observed values but also the residuals of the embedded physical equations, ensuring that the model is in line with the representation of the physical information while improving the accuracy and robustness of the prediction.
However, directly applying physical constraints to diffusion models is challenging. Real-world radio environments are too complex to be described by a single, strict equation, and diffusion models predict noise, making direct physical loss application impractical. To address this, we designed a dual U-Net architecture. One U-Net (U-Net 1) is constrained by an approximate two-dimensional Helmholtz equation to provide physical representation guidance. In parallel, a second U-Net (U-Net 2) (Ronneberger et al., 2015) refines these features through a diffusion process, generating high-resolution radio maps. By leveraging the Helmholtz equation to inform the learning process, we enhance the model’s ability to generate accurate predictions and improve its generalization in scenarios with sparse or unevenly distributed data.
@@
This physical information representation alignment and machine learning principles enables PhyRMDM to accurately capture complex spatial signal characteristics (Raissi et al., 2019a). In experiments conducted under the Static RM (SRM) setting, our model achieved an NMSE of 0.0031 and an RMSE of 0.0125, outperforming state-of-the-art (SOTA) approaches. Extensive experimental validation further highlights the model’s superior performance in reconstructing radio maps across various challenging scenarios, establishing a novel paradigm that combines physics-driven methodologies with data-driven techniques to address real-world signal propagation challenges (Li et al., 2020).
@@
  • We propose the PhyRMDM, a physics-informed representation consistency framework that integrates the Helmholtz equation and diffusion processes within a dual U-Net architecture. This design ensures physical consistency and significantly enhances the accuracy of radio map reconstruction in sparse and complex propagation scenarios.
@@
  • By aligning representation of Physics-Informed Neural Networks (PINNs), PhyRMDM embeds physical information representation alignment into the learning process, enabling precise modeling of spatial signal characteristics, such as signal strength and path loss, while improving generalization across diverse environments.
@@
  • Extensive experimental validation demonstrates that PhyRMDM achieves state-of-the-art performance, including an NMSE of 0.0031 and RMSE of 0.0125 under the Static RM (SRM) setting, highlighting its robustness and scalability for real-world signal propagation challenges.
2. Related Work
2.1. Radio Map Estimation
@@
Radio map estimation is essential for applications such as network planning and spectrum management. Traditional methods include model-based approaches, which utilize established signal propagation models like the log-distance path loss (LDPL) model for Wi-Fi radio map reconstruction (Jung et al., 2011). Interpolation techniques, such as thin-plate splines (Keller and Borkowski, 2019), have also been employed to achieve spatially continuous estimations. However, these methods often assume ideal conditions and may not accurately capture complex environmental factors like shadowing and multipath effects, leading to limitations in heterogeneous settings.
2.2. Diffusion Models
Diffusion models can capture complex data distributions and have been applied to radio map estimation. RadioDiff (Nichol and Dhariwal, 2021) enables sampling-free dynamic radio map construction, while RM-Gen (Luo et al., 2024) uses conditional denoising diffusion to synthesize maps from minimal data. These methods reduce reliance on extensive measurements and costly simulations.
2.3. Physics-Informed Neural Networks (PINNs)
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019b) incorporate physical laws, typically expressed as partial differential equations (PDEs) (Sloan et al., 2012), into neural network training to ensure predictions comply with known physics, improving interpretability and generalization in sparsedata scenarios. For radio map estimation, embedding the Helmholtz equation—describing wave propagation—into the loss function enhances accuracy and robustness under challenging measurement conditions.
PINNs have been applied across diverse scientific and engineering domains (Hu and McDaniel, 2023; Eghtesad et al., 2024), such as solving the Navier–Stokes equations in fluid mechanics, incorporating constitutive laws in materials science, improving climate forecasts via conservation laws, and modeling electromagnetic wave propagation. These applications demonstrate that physical constraints reduce overfitting and boost interpretability, especially with limited or noisy data.
Unlike conventional PINNs that enforce PDE residuals only at the output, we adopt a Representation Alignment strategy that aligns latent features with high-quality priors, ensuring both physical consistency and semantic discriminability.
3. Method
@@
Figure 2. An illustration of the PhyRMDM framework. (a) shows the pipeline overview, combining a condition model and a diffusion model for radio map reconstruction. Key components include (b) a physics-constrained feature alignment module, which integrates physical priors with learned representations, and (c) Radio Frequency Spatial Attention (RF-SA), which aligns spatial and frequency features to enhance accuracy. The framework leverages anchor conditions and representation alignment to iteratively refine features, ensuring accurate predictions in complex propagation scenarios.
@@
The radio map estimation task aims to predict the spatial distribution of wireless signal strength within a given area. Wireless signal propagation follows physical laws, such as electromagnetic wave equations and path loss models (Jiang et al., 2024). Incorporating these physical constraints into learning-based models can significantly improve the accuracy and reliability of predictions. Here is a complementary approach we have developed : aligning learned representations with domain-specific priors to improve model efficiency and robustness.
@@
As illustrated in Figure 2, we propose a diffusion-based probabilistic framework that integrates a dual UNet architecture with lightweight physical constraint modules. The framework emphasizes representation alignment between multi-scale features and domain-specific signal characteristics, structured in two stages:
@@
  • Conditional Generation Model (UNet 1):In this stage, input features are extracted and a Physics-Informed Neural Network (PINN) module enforces physical constraints to achieve physical information representational alignment.
@@
3.1. Problem Formulation
Reconstructing the radio map involves modeling the spatial distribution of wireless signals, which can be formulated using the Helmholtz equation (Juraev et al., 2024). This equation provides a physical foundation for capturing signal propagation dynamics:
(1) $\nabla^{2}u(x,y)+k^{2}(x,y)u(x,y)=-f(x,y),\quad(x,y)\in\Omega.$
@@ 3.1. Problem Formulation
(2) $\Delta_{h}u(i,j)$ $=\frac{u(i+1,j)+u(i-1,j)+u(i,j+1)+u(i,j-1)}{h^{2}}$
$\qquad-\frac{4u(i,j)}{h^{2}},$
@@
(4) $\mathcal{L}_{\text{PDE}}=\frac{1}{N_{\text{int}}}\sum_{(i,j)\in\Omega_{\text{int}}}[r(i,j)]^{2},$
@@
(5) $\mathcal{L}_{\text{BC}}=\frac{1}{N_{\text{bc}}}\sum_{(i,j)\in\partial\Omega}[u(i,j)-u_{\text{BC}}(i,j)]^{2},$
@@
(6) $\mathcal{L}_{\text{Source}}=\frac{1}{N_{\text{src}}}\sum_{(i,j)\in\Omega_{\text{src}}}[u(i,j)-u_{\text{src}}(i,j)]^{2},$
@@
(7) $\mathcal{L}_{\text{PINN}}=\lambda_{\text{PDE}}\mathcal{L}_{\text{PDE}}+\lambda_{\text{BC}}\mathcal{L}_{\text{BC}}+\lambda_{\text{Source}}\mathcal{L}_{\text{Source}},$
@@
3.2. Model Architecture
Directly applying physical constraints to diffusion models is challenging, as they predict noise rather than signal strength, and complex radio environments are difficult to model with a single equation. To address this, our model employs a dual U-Net architecture. One U-Net (U-Net 1) is constrained by an approximate two-dimensional Helmholtz equation to provide physical representation guidance. In parallel, a second U-Net (U-Net 2) refines these features through a diffusion process to generate the final radio map (Kawar et al., 2022). To enhance feature transfer between the two U-Nets, we propose two innovative mechanisms: the Anchor Condition and RF-SA.
3.2.1. UNet 1: Conditional Generation Model with PINN Alignment